# HD-PiSSA: High-Rank Distributed Orthogonal Adaptation

### v1

@@ Abstract
Existing parameter-efficient fine-tuning (PEFT) methods for large language models (LLMs), such as LoRA and PiSSA, constrain model updates to low-rank subspaces, limiting their expressiveness and leading to suboptimal performance on complex tasks. To address this, we introduce High-rank Distributed PiSSA (HD-PiSSA), a distributed PEFT approach that initializes orthogonal adapters across different devices and aggregates their delta updates collectively on $W$ for fine-tuning. Unlike Data Parallel LoRA or PiSSA, which maintain identical adapters across all devices, HD-PiSSA assigns different principal components of the pre-trained weights to each GPU, significantly expanding the range of update directions. This results in over 16$\times$ higher effective updated ranks than data-parallel LoRA or PiSSA when fine-tuning on 8 GPUs with the same per-device adapter rank. Empirically, we evaluate HD-PiSSA across various challenging downstream tasks, including mathematics, code generation, and multi-task learning. In the multi-task setting, HD-PiSSA achieves average gains of 10.0 absolute points (14.63%) over LoRA and 4.98 points (6.60%) over PiSSA across 12 benchmarks, demonstrating its benefits from the extra optimization flexibility.
@@ 5 Experiments
Table 1: Accuracy comparison of LLaMA-2-7B, LLaMA-3.1-8B, and Mistral-7B-v0.1 with LoRA, PiSSA, and HD-PiSSA on mathematics and code generation datasets.
Model Method Params GSM8K MATH HumanEval MBPP Avg.
@@ 5.2 Performance for Multi-Task Fine-Tuning
We further evaluate HD-PiSSA on a widely adopted multi-task commonsense reasoning setting Liu et al. (2024), comparing it against high-rank PEFT methods such as MoRA, HiRA, and DoRA. Training is conducted on eight standard datasets (BoolQ, PIQA, SIQA, HellaSwag, WinoGrande, ARC-c, ARC-e, and OBQA), and evaluated on their respective test sets. To ensure fairness, we match the number of trainable parameters by tuning only the $q\_proj,k\_proj,v\_proj$ and $up,down$ matrices for three epochs. We report results from the final model without validation-based checkpointing. As shown in Table 2, HD-PiSSA achieves a 2.0% average gain over the previous state-of-the-art (HiRA), and consistently outperforms other high-rank approaches.
@@ 5.5 More Ablation Study
We future conducted ablation studies to verify the effectiveness of Direct Weight Update (DWU) and Orthogonal Adapter Initialization (OAI), as well as the impact of different mute scalar values. Detailed results are provided in the Appendix C and  D.
@@ 6 Analysis of Effective Update Rank
Figure 7: Comparison of LoRA(r=128), PiSSA(r=128) and HD-PiSSA(r=128)’s updated ranks. The singular values of HD-PiSSA remain 1 to 2 orders of larger than LoRA and PiSSA beyond $2Kr$.
@@ 8 Limitations
HD-PiSSA requires multi-GPU distributed setups to fully realize its benefits, which may restrict its applicability in resource-constrained environments. Additionally, while demonstrating strong results on various tasks, its generalization to broader domains and modalities remains to be validated. We also plan to conduct more rigorous theoretical analysis on the effective update rank as part of future work.
@@ A.4 Training Setting in Main Experiments
Precision float32
@@ Appendix B Multi-task Fine-Tuning Dataset
Appendix C Ablation Study on Direct Weight Update (DWU) and Orthogonal Adapter Initialization (OAI)
To further understand the design choices in HD-PiSSA, we conducted an ablation study on two key components: Direct Weight Update (DWU) and Orthogonal Adapters Initialization (OAI). Specifically, we included:
  • LoRA init. (with DWU): A variant using standard LoRA initialization with the Direct Weight Update technique.
  • Top-Rank HD-PiSSA (w/o OAI): HD-PiSSA without orthogonal adapters, using PiSSA adapter instead.
  • HD-PiSSA: The full version with both DWU and OAI.
All variants were fine-tuned on the code dataset for 3 epochs. Evaluation was conducted on HumanEval and MBPP (abbreviated as H&M in the table) every 500 steps for detailed training-time insights. The total batch size and training configuration were kept consistent across methods.
Components / Steps Avg. Loss 500 1000 1500 2000 Final (2323)
Analyzing the results in Table 7, we observe that HD-PiSSA’s superior performance stems from the synergy of both DWU and OAI. First, the use of DWU clearly provides a measurable gain, as seen in the improved performance of LoRA+DWU compared to standard baselines. This aligns with findings in prior work such as AdaLoRA. However, even with DWU, both LoRA and PiSSA-based variants lag behind HD-PiSSA in terms of average loss and final evaluation metrics.
These results confirm that Orthogonal Adapters Initialization (OAI) plays a crucial role in HD-PiSSA’s performance. The full version consistently outperforms all ablated variants, highlighting the importance of orthogonal modularity for stable and effective multi-component training.
Appendix D Performances with Different Mute Scalars
Figure 8: Performances on GSM8K with different mute scalars.
In this section, we evaluate the impact of different mute scalar values on performance. We tested several mute scalar values, including $1\times 10^{-4}$, $1\times 10^{-8}$, $1\times 10^{-16}$, $1\times 10^{-2}$, and $1\times 10^{-32}$. Specifically, we finetune LLaMA-2-7B with these mute scalars on MetaMathQA for one epoch, then evaluate the performance on GSM8K test set. As shown in Fig. 8, performance remained consistent for mute scalar values of $1\times 10^{-4}$, $1\times 10^{-8}$, and $1\times 10^{-16}$. However, when the mute scalar was set to $1\times 10^{-2}$ or $1\times 10^{-32}$, performance decreased significantly.
The likely cause of this behavior is that if the mute scalar is relative large, the adapter’s contributions during forward propagation are not fully eliminated, which affects the output $y$. On the other hand, setting the scalar too small might result in reaching the lower limit of float32 precision, causing inaccuracies in gradient updates.

### v3

@@ Abstract
[…]
@@ 5 Experiments
Table 1: Accuracy comparison of LLaMA-2-7B, LLaMA-3.1-8B, and Mistral-7B-v0.1 with LoRA, PiSSA, and HD-PiSSA on mathematics and code generation datasets. “Params.” denotes the per-GPU optimizer parameters, which remain the same across LoRA, PiSSA, and HD-PiSSA.
Model Method Params. GSM8K MATH HumanEval MBPP Avg.
@@ 5.2 Performance for Multi-Task Fine-Tuning
We further evaluate HD-PiSSA on a widely adopted multi-task commonsense reasoning setting Liu et al. (2024), comparing it against high-rank PEFT methods such as MoRA, HiRA, and DoRA. Training is conducted on eight standard datasets (BoolQ, PIQA, SIQA, HellaSwag, WinoGrande, ARC-c, ARC-e, and OBQA), and evaluated on their respective test sets. To ensure fairness, we match the number of trainable parameters by tuning only the q_proj,k_proj,v_projq\_proj,k\_proj,v\_proj and $up,down$ matrices for three epochs. We report results from the final model without validation-based checkpointing. As shown in Table 2, HD-PiSSA achieves a 2.0% average gain over the previous state-of-the-art (HiRA), and consistently outperforms other high-rank approaches.
6 Ablation Study
@@ 6.2 Experiments on Various Ranks
6.3 Performances with Different Mute Scalars
Figure 7: Performances on GSM8K with different mute scalars.
We also evaluate the impact of different mute scalar values on performance. We tested several mute scalar values, including $\times 10^{-4}, 1×1$$\times 10^{-8}$, $\times 10^{-16}, 1×1$$\times 10^{-2}$, and $1\times 10^{-32}$. Specifically, we finetune LLaMA-2-7B with these mute scalars on MetaMathQA for one epoch, then evaluate the performance on the GSM8K test set. As shown in Fig. 7, performance remained consistent for mute scalar values of $\times 10^{-4}, 1×1$$\times 10^{-8}$, and $1\times 10^{-16}$. However, when the mute scalar was set to $1\times 10^{-2}$ or $1\times 10^{-32}$, performance decreased significantly.
The likely cause of this behavior is that if the mute scalar is relatively large, the adapter’s contributions during forward propagation are not fully eliminated, which affects the output $y$. On the other hand, setting the scalar too small might result in reaching the lower limit of float32 precision, causing inaccuracies in gradient updates.
@@ 6.4 More Ablation Study
We conducted future ablation studies to verify the effectiveness of Direct Weight Update (DWU) and Orthogonal Adapter Initialization (OAI). Specifically, we included: (1) LoRA init. (with DWU): A variant using standard LoRA initialization with the Direct Weight Update technique. (2) Top-Rank HD-PiSSA (w/o OAI): HD-PiSSA without orthogonal adapters, using PiSSA adapter instead. All variants were fine-tuned on the code dataset for 3 epochs. Evaluation was conducted on HumanEval and MBPP (abbreviated as H&M in the table) every 500 steps for detailed training-time insights. The total batch size and training configuration were kept consistent across methods.
Components / Steps Avg. Loss 500 1000 1500 2000 Final (2323)
Analyzing the results in Table 3, we conclude that HD-PiSSA’s superior performance is due to the synergy of DWU and OAI. First, the use of DWU clearly provides a measurable gain, as seen in the improved performance of LoRA+DWU compared to standard baselines. This aligns with findings in prior work such as AdaLoRA. However, even with DWU, both LoRA and PiSSA-based variants lag behind HD-PiSSA in terms of average loss and final evaluation metrics.
We also test HD-PiSSA’s training overhead compared with PiSSA, and found that HD-PiSSA’s time and memory costs slightly increase. Details in the Appendix C
@@ 7 Analysis of Effective Update Rank
Figure 8: Comparison of singular value magnitudes for the update matrices of LoRA ($r=128$), PiSSA ($r=128$), and HD-PiSSA ($r=128$). The sub-figure in the top-right corner provides a magnified view of the singular values for indices beyond $2Kr$. This highlights that the singular values for HD-PiSSA remain 1 to 2 orders of magnitude larger than those of LoRA and PiSSA in this high-rank region, empirically supporting that HD-PiSSA achieves update rank beyond $2Kr$.
@@ 9 Limitations
HD-PiSSA requires multi-GPU distributed setups to fully realize its benefits, which may restrict its applicability in resource-constrained environments. Besides, because HD-PiSSA aggregates all the gradient updates on the Pretrained Matrix W, its performance decreases when the base model is loaded at lower precision. Additionally, while demonstrating strong results on various tasks, its generalization to broader domains and modalities remains to be validated. We also plan to conduct more rigorous theoretical analysis on the effective update rank as part of future work.
@@ A.4 Training Setting in Main Experiments
Details are presented in Table 5. Results are obtained from a single time running. Our evaluation codebase follows PiSSA’s open-source evaluation codebase.
@@
Base Model Precision float32
@@ Appendix B Multi-task Fine-Tuning Dataset
Appendix C Speed Comparison
Table 8: Training overhead comparison between PiSSA and HD-PiSSA on LLaMA-2-7B using 8 GPUs. Both methods are run for 100 steps on the MetaMath dataset without advanced memory or acceleration optimizations.
Metric PiSSA HD-PiSSA
Peak Mem. 47,290 MiB 49,176 MiB
Time (100 steps) 1808.31 s 1995.8 s
To evaluate the training overhead of our proposed HD-PiSSA, we conducted a comparative analysis against PiSSA. The experiment was performed on the MetaMath dataset using a LLaMA-2-7B model, distributed across 8 GPUs with 16 ranks per GPU, and using bf16 precision without any advanced memory and acceleration optimizations (our implementation of parallelization and communication).
As detailed in Table 8, HD-PiSSA exhibits a slight increase in both peak memory usage per GPU (49,176 MiB vs. 47,290 MiB) and total training time for 100 steps (1995.8s vs. 1808.31s). We attribute this marginal overhead to the additional communication and computation required for updating the full-parameter matrix $W$. For instance, each GPU is required to gather the previous values and gradients of matrices $A$ and $B$ from all other GPUs, followed by computing the corresponding update to $W$. Nevertheless, we consider this overhead to be acceptable given the substantial gains in performance demonstrated by HD-PiSSA.